OpenStudy (anonymous):

Simplify the complex algebraic expression: (12x^3)/(7y^4)/(3x^5)/(2y)

OpenStudy (mathlegend):

Are they two fractions?

OpenStudy (anonymous):

Yes.

OpenStudy (mathlegend):

So upon having division of fractions... we never really divide. We multiply instead. So in order to do this... we need to flip the 2nd fraction. (Reciprocal)

OpenStudy (mathlegend):

\[\frac{ 12x ^{3} }{ 7y ^{4} }*\frac{ 2y }{ 3x ^{5} }\]

OpenStudy (mathlegend):

So multiply straight across please and tell me what you get.

OpenStudy (anonymous):

(24x^3y)/(21x^5y4)?

OpenStudy (mathlegend):

Yes!

OpenStudy (mathlegend):

Now, can't we simplify? Lets start with the whole numbers out front...

OpenStudy (mathlegend):

\[\frac{ 24 }{ 21 }\]

OpenStudy (mathlegend):

What does that reduce to?

OpenStudy (anonymous):

8/7..

OpenStudy (mathlegend):

Good... so now we have...

OpenStudy (mathlegend):

\[\frac{ x ^{3}y }{ x ^{5}y ^{4} }\]

OpenStudy (mathlegend):

Now, lets do the x's first... we know that 5-3 = 2 right?

OpenStudy (anonymous):

(8/7x^3y^3

OpenStudy (mathlegend):

Since the 5 is the bigger number it determines where the x^2 will end up... 5 is on the bottom so that is where the x^2 will go

OpenStudy (mathlegend):

Wow, @Katelynmorris you are good, just minor error... it should be x^3

OpenStudy (mathlegend):

x^2

OpenStudy (mathlegend):

sorry!

OpenStudy (anonymous):

Thank you! :)